Saved Help Save &Exit Su You buy an 7-year $1,000 par value bond today that has a 540% yield and a 540% annual payment coupon. In 1 year prom sed yields have risen to 6.40%. Your 1year holding period return was-. Multiple Choice 1.08% -486% -272% 0.54%
Answer #1

Answer:

Correct answer is:

0.54%

Explanation:

At the time of purchase:

Annual coupon rate = 5.40%

Yield = 5.40%

Hence purchase price = Par value = $1000

Period = 7 years

Annual coupon amount = 1000 * 5.40% = $54

In 1 year promised yield has risen to 6.40%

To calculate price after one year we use PV function of excel:

PV (rate, nper, pmt, fv, type)

= PV (6.40%, 6, -54, -1000,0)

= $951.44

Price after one year = $951.44

As such 1 year holding period return = (Price after one year - Purchase price + Annual coupon) / Purchase price

= (951.44 - 1000 + 54) / 1000

= 0.54%

As such option D is correct and other options A, B and C are incorrect.